# ND-SDF: Learning Normal Deflection Fields for High-Fidelity Indoor Reconstruction

### v1

@@ Abstract
Neural implicit reconstruction via volume rendering has demonstrated its effectiveness in recovering dense 3D surfaces. However, it is non-trivial to simultaneously recover meticulous geometry and preserve smoothness across regions with differing characteristics. To address this issue, previous methods typically employ geometric priors, which are often constrained by the performance of the prior models. In this paper, we propose ND-SDF, which learns a Normal Ddeflection field to represent the angular deviation between the scene normal and the prior normal. Unlike previous methods that uniformly apply geometric priors on all samples, introducing significant bias in accuracy, our proposed normal deflection field dynamically learns and adapts the utilization of samples based on their specific characteristics, thereby improving both the accuracy and effectiveness of the model. Our method not only obtains smooth weakly textured regions such as walls and floors but also preserves the geometric details of complex structures. In addition, we introduce a novel ray sampling strategy based on the deflection angle to facilitate the unbiased rendering process, which significantly improves the quality and accuracy of intricate surfaces, especially on thin structures. Consistent improvements on various challenging datasets demonstrate the superiority of our method.
@@ 1 Introduction
Recently, coordinate-based networks Mildenhall et al. (2021); Barron et al. (2022); Müller et al. (2022); Martin-Brualla et al. (2021); Zhang et al. (2020); Ye et al. (2023a); Huang et al. (2024); Ming et al. (2022) are drawing increasing attention, due to their remarkable performance in the task of novel view synthesis. Inspired by the implicit representation of the scene, many subsequent works introduced SDF Park et al. (2019) or occupancy Li et al. (2022) to parameterize the 3D geometry.
However, recovering high-fidelity surfaces remains challenging, as relying solely on color images for supervision often results in an underconstrained problem, especially for regions like textureless walls and ceilings. Previous methods Yu et al. (2022b); Wang et al. (2022) have attempted to mitigate this issue by incorporating auxiliary supervision. For instance, MonoSDF employs monocular cues from a pretrained model as pseudo-ground truth to supervise the model, which partially alleviates the problem in textureless areas. However, the large errors due to domain gaps in monocular cues and the inconsistencies in view-dependent prior guidance often lead to visible detail loss and erroneous surfaces, as depicted in Figure 3 of the experimental section.
Noting that prior models are highly accurate in simpler regions like floors and walls but struggle with complex geometries, we introduce ND-SDF for high-fidelity indoor reconstruction (Figure 1). Our core innovation is the construction of a deflection field that adaptively learns the geometric deviations between the actual scene geometry and the prior geometry derived from normal priors. The inherent deviation is defined as the angular difference between the true scene normals and the prior normals. By aligning the deflected scene normals with the normal cues, our model can adaptively learn the deviation, which encourages accurate recovery of intricate structures maintaining fine details without being misled by erroneous priors.
In addition, we propose a novel adaptive deflection angle prior loss that leverages prior normals for differential supervision of high and low-frequency areas, achieving an optimal balance between smoothness and detail. Our observations indicate that large angle deviations are primarily concentrated in thin and fine-grained structures. Building on this insight, we introduce deflection-angle guided optimization to proactively facilitate the recovery of detailed structures. Furthermore, we address a significant bias issue and propose deflection angle-guided unbiased rendering to improve the reconstruction of small or thin structures, as shown in Figure 1. In summary, we present the following contributions:
@@ 2 Related Work
Due to the complex layouts of indoor scenes, additional auxiliary data are required for reasonable reconstruction. Manhattan-SDF Guo et al. (2022) employs the Manhattan assumption and semantic priors to jointly regularize textureless regions such as walls and floors. HelixSurf Liang et al. (2023) achieves intertwined regularization iteratively by combining neural implicit surface learning with PatchMatch-based multi-view stereo (MVS) Barnes et al. (2009) robustly and efficiently. Other works Yu et al. (2022b); Wang et al. (2022) propose utilizing monocular priors from pretrained models to achieve smooth and complete reconstruction. However, directly applying normal or depth priors Yu et al. (2022b) may lead to undesirable reconstruction results due to the unreliable nature of these priors. NeuRis Wang et al. (2022) filters out unreliable priors based on the assumption that areas with rich 2D visual features are more error-prone. H2O-SDF Park et al. (2024) simply utilizes prior uncertainty to re-weight the normal prior loss. Nevertheless, both of these methods exhibit weak generalization capabilities. For example, the visual feature hypothesis Wang et al. (2022) struggles to differentiate between rich-textured planar areas. Similarly, prior-guided re-weighting Park et al. (2024) is further constrained by the domain gap between the prior model and the scene. On the other hand, DebSDF Xiao et al. (2024) effectively designs an uncertainty field based on a probabilistic model to guide the loss function, leading to more robust and accurate reconstructions. In contrast to methods with weaker generalization capabilities and those relying on probabilistic models, our proposed method, ND-SDF, learns a Normal Deflection field that enables the dynamic adaptation of priors based on their characteristics. This field directly models geometrically meaningful SO(3) residuals, resulting in highly detailed surface reconstructions while maintaining smoothness and robustness.
Figure 2: Overview of our method. We utilize multi-resolution hash grids $\gamma_{L}$ as scene representation. The core of ND-SDF is the normal deflection field. We represent deflection with quaternions, which are predicted by the deflection network (denoted as $f_{d}$). We align the deflected normals with the prior normals to learn the deviation between the scene and the priors. To distinctly supervise high and low-frequency areas, we employ an adaptive deflection angle prior loss, ensuring both smoothness and detail. Furthermore, we utilize the deflection angle $\Delta\theta$ to distinguish complex structures, enabling angle-guided sampling and color loss to facilitate intricate surface details. Lastly, we combine the unbiased rendering method Zhang et al. (2023) to ensure the generation of extremely thin structures indoors.
@@ 3 Method
The primary objective of our Normal Deviation Signed Distance Function (ND-SDF) is to reconstruct dense surfaces from calibrated multi-view images, with a primary focus on indoor scenes with established priors. To achieve this, we introduce a novel component known as the normal deflection field. This field is designed to quantify and learn the deviations between actual scene geometries and their corresponding normal priors. By integrating this field, our method can adaptively supervise both high and low-frequency regions in the scene. This adaptive supervision is crucial for preserving fine details while ensuring overall surface smoothness.
@@ 3.1 Preliminaries
Volume Rendering NeRF assumes that a ray $\textbf{r}(t)=\textbf{o}+t\textbf{v}$ is emitted from viewpoint o in direction v, where $t$ denotes the distance from the viewpoint. $N$ points are sampled on the ray, i.e. $\textbf{x}_{i}=\textbf{o}+t_{i}\textbf{v},i\in\{1,2...N\}$. Given each point’s volume density $\sigma_{i}$ and color $\textbf{c}_{i}$, the color of the ray can be synthesized using volume rendering techniques as:
@@
where $\alpha$ is the opacity of $i$-th ray segment, $T$ denotes the transmittance rate, $\delta_{i}=t_{i}-t_{i-1}$ denotes distance between adjacent samples.
A color loss supervising the rendered color and ground-truth color is utilized to optimize the network:
@@
where $\mathcal{R}$ is the sampled ray set.
SDF-induced Volume Rendering The rise of neural surface reconstruction is attributed to methods like VolSDF and Neus, which propose specific SDF to volume density representation, allowing optimization with volume rendering. The final surface, denoted as $S$, is equivalent to the zero level-set of the implicit distance field, i.e., $S=\{\textbf{x}\in\mathbb{R}^{3}|s(\textbf{x})=0\}$, where $s(\textbf{x})$ is the SDF value.
We follow VolSDF, employing the Laplace Cumulative Distribution Function(CDF) to model the relationship between SDF and volumetric density:
@@
where $\Psi_{\beta}$ denotes the Laplace CDF, $\beta$ denotes the variance. The depth $\hat{D}(\textbf{r})$ and normal $\hat{\textbf{N}}(\textbf{r})$ of the surface intersection point is synthesized using volume rendering technology:
$\hat{D}(\textbf{r})=\sum_{i=1}^{N}T_{i}\alpha_{i}t_{i},\hat{\textbf{N}}(\textbf{r})=\sum_{i=1}^{N}T_{i}\alpha_{i}\textbf{n}_{i},$ (4)
where $\textbf{n}_{i}$ is the analytical gradient of the SDF network. Following MonoSDF Yu et al. (2022b), we utilize monocular depth and normal generated by pretrained model Eftekhar et al. (2021) to supervise the rendered depth and normal:
@@ 3.2 Normal deflection field
A significant challenge in indoor 3D reconstruction is achieving a balance between the smoothness of surfaces and the intricacy of complex structures. Traditional approaches relying solely on photometric loss have proven inadequate for accurately capturing smooth areas, particularly in texture-deficient regions. These methods often necessitate auxiliary data, such as normal priors, to enhance the reconstruction quality in textureless areas. However, the uniform application of normal priors across different scene types can impair the recovery of complex structures. This is primarily due to the variable reliability of these priors in diverse regions, where they may not accurately reflect the underlying geometry. To address this issue, we propose the development of a Normal Deflection field. This field is designed to dynamically represent the deviation between the actual scene normals and the provided normal priors. By doing so, it effectively circumvents the potential misguidance from inconsistent priors, thereby enabling a more reliable and nuanced reconstruction of both smooth and complex indoor structures.
We choose quaternion as the deflection form, which is a lightweight rotation representation. The quaternion can be parameterized by a single MLP $f_{d}$:
@@
where $\textbf{q}_{i}=(\textbf{q}_{i}^{0},\textbf{q}_{i}^{1},\textbf{q}_{i}^{2},\textbf{q}_{i}^{3})$ is the deflection quaternion(normalized default) of sampled $\textbf{x}_{i}$. Quaternion at the surface intersection point is synthesized using volume rendering technology, similar to how NeRF synthesizes colors:
@@ 3.3 Adaptive deflection angle prior loss
The learned deviation is defined as the deflection angle($\Delta\theta$) of the rendered normal:
@@
where $\Delta\theta\in[0,\pi]$, Empirically, the deviation between the scene and priors increases as the structure becomes more complex. It leads to an intuitive conclusion that the deflection angle is small in smooth areas and large in high-frequency areas. The adaptive deflection angle normal prior loss is proposed to dynamically adjust the utilization of differing priors based on their characteristics:
$\begin{array}[]{cc}\mathcal{L}^{ad}_{normal}=&\sum_{\textbf{r}\in\mathcal{R}}{g^{d}(\Delta\theta)\mathcal{L}^{d}_{normal}(\hat{\textbf{N}}^{d}(\textbf{r}),{\textbf{N}}(\textbf{r}))}+\\ &g(\Delta\theta)\mathcal{L}_{normal}(\hat{\textbf{N}}(\textbf{r}),{\textbf{N}}(\textbf{r}))\end{array},$ (12)
where $g^{d}$ and $g$ are modulation functions that adjust the weight of the deflected and original normal loss term based on the deflection angle. We similarly define an adaptive depth prior loss:
@@
As $\Delta\theta$ increases, $g^{d}(\Delta\theta)$ increases, and $g(\theta)$ decreases, since we should less apply priors in complex areas indicated by large deflection angles. The combined adaptive normal and depth prior loss is referred to as the adaptive deflection angle prior loss(see appendix for detailed definitions).
@@ 3.4 Deflection angle guided optimization
Through the utilization of the proposed adaptive prior loss, we dynamically adjust the utilization of various priors, significantly enhancing the quality of reconstruction. However, this alone is insufficient for generating more complex structures. Fundamentally, merely learning the deviation does not endow the capability to reconstruct additional details. Recognizing that large deflection angles indicate complex areas, we introduce three deflection angle guided optimization methods designed to facilitate the recovery of thinner and more fine-grained structures.
@@
Textureless areas, such as walls, constitute a significant portion of indoor scenes. These areas converge rapidly by utilizing priors. To prevent continuous oversampling of well-learned smooth regions, we proactively sample more rays in complex areas to capture finer details. The sampling process is guided by the deflection angles (see supplementary for details).
@@
To further promote the recovery of details, we impose additional photometric loss on complex areas indicated by the deflection angles. We re-weight the original color loss guided by the deflection angle, resulting in the re-weighted color loss:
@@
where ${\Delta\theta}(\textbf{r})$ denotes the deflection angle of sampled ray r, and $w_{color}$ is the re-weight function guided by ${\Delta\theta}$.
@@
Accurately reconstructing thin structures, such as chair legs, remains challenging. This limitation arises from inherent bias issues in SDF-induced volume rendering, manifesting in two ways: (1) According to TUVR Zhang et al. (2023), the derivative of the rendering weight $\frac{\partial w(t)}{\partial t}$ is influenced by various angle differences between ray directions and scene normals. This results in a non-maximum value of $w$ at the surface intersection point, reducing surface quality. (2) For rays passing close to the object, the Laplace CDF assigns high density to points near the surface of the ray, leading to incorrect rendering depths and normals. This issue causes thin structures to gradually disappear during training. Therefore, it is crucial to apply an unbiased rendering method during reconstruction.
Following TUVR Zhang et al. (2023), we transform SDF to density using an unbiased function:
$\sigma(\textbf{r}({t_{i}}))=\alpha\Psi_{\beta}\left(\frac{-f_{g}(\textbf{r}({t_{i}}))}{\left|f_{g}^{\prime}(\textbf{r}({t_{i}}))\right|}\right),$ (15)
where $\textbf{r}(t_{i})=\textbf{o}+t_{i}\textbf{v}$ represents a sampled point on the ray, and $f_{g}$ is the geometry network, which predicts the SDF value. This transformation effectively alleviates bias issues. However, we observe that simply applying it results in the surface failing to converge. Hence, we only partially apply it to thin structures indicated by the learned deflection angles.

### v2

@@ Abstract
Neural implicit reconstruction via volume rendering has demonstrated its effectiveness in recovering dense 3D surfaces. However, it is non-trivial to simultaneously recover meticulous geometry and preserve smoothness across regions with differing characteristics. To address this issue, previous methods typically employ geometric priors, which are often constrained by the performance of the prior models. In this paper, we propose ND-SDF, which learns a Normal Deflection field to represent the angular deviation between the scene normal and the prior normal. Unlike previous methods that uniformly apply geometric priors on all samples, introducing significant bias in accuracy, our proposed normal deflection field dynamically learns and adapts the utilization of samples based on their specific characteristics, thereby improving both the accuracy and effectiveness of the model. Our method not only obtains smooth weakly textured regions such as walls and floors but also preserves the geometric details of complex structures. In addition, we introduce a novel ray sampling strategy based on the deflection angle to facilitate the unbiased rendering process, which significantly improves the quality and accuracy of intricate surfaces, especially on thin structures. Consistent improvements on various challenging datasets demonstrate the superiority of our method.
@@ 1 Introduction
Recently, coordinate-based networks Mildenhall et al. (2021); Barron et al. (2022); Müller et al. (2022); Martin-Brualla et al. (2021); Zhang et al. (2020); Ye et al. (2023a); Huang et al. (2024); Ming et al. (2022) are drawing increasing attention, due to their remarkable performance in the task of novel view synthesis. Inspired by implicit representations of the scene, many subsequent works have introduced Signed Distance Functions (SDF) Park et al. (2019) or occupancy Li et al. (2022) to parameterize the 3D geometry. Notably, some of these approaches Yariv et al. (2021); Wang et al. (2021) outperform other methods in recovering dense surfaces by representing scene geometry with implicit distance fields and integrating them with differentiable rendering frameworks to enhance surface reconstruction.
However, recovering high-fidelity surfaces remains challenging, as relying solely on color images for supervision often results in an underconstrained problem, especially for regions like textureless walls and ceilings. Previous methods Yu et al. (2022b); Wang et al. (2022) have attempted to mitigate this issue by incorporating auxiliary supervision. For instance, MonoSDF employs monocular cues from a pretrained model as pseudo-ground truth to supervise the reconstruction process, which alleviates the issue in textureless regions. However, the large errors due to domain gaps in monocular cues and the inconsistencies in view-dependent prior guidance often lead to visible detail loss and erroneous surfaces Yu et al. (2022b), as depicted in Figure 4 of the experimental section.
Noting that prior models are highly accurate in simpler regions like floors and walls but struggle with complex geometries. Based on this insight, we introduce ND-SDF for high-fidelity indoor reconstruction (Figure 1). Our core innovation is the construction of a deflection field that adaptively learns the geometric deviations between the actual scene geometry and the prior geometry derived from normal priors. The inherent deviation is defined as the angular difference between the true scene normals and the prior normals. By aligning the deflected scene normals with the normal cues, our model can adaptively learn the deviation, which encourages accurate recovery of intricate structures maintaining fine details without being misled by erroneous priors.
In addition, we propose a novel adaptive deflection angle prior loss that leverages prior normals for differential supervision of high and low-frequency areas, achieving an optimal balance between smoothness and detail. Our observations indicate that large angle deviations are primarily concentrated in thin and fine-grained structures. Building on this insight, we introduce deflection-angle guided optimization to proactively facilitate the recovery of detailed structures. Furthermore, we address a significant bias issue and propose deflection angle guided unbiased rendering to improve the reconstruction of small or thin structures, as shown in Figure 1. In summary, we present the following contributions:
@@ 2 Related Work
Due to the complex layouts of indoor scenes, additional auxiliary data are required for reasonable reconstruction. Manhattan-SDF Guo et al. (2022) employs the Manhattan assumption and semantic priors to jointly regularize textureless regions such as walls and floors. HelixSurf Liang et al. (2023) achieves intertwined regularization iteratively by combining neural implicit surface learning with PatchMatch-based multi-view stereo (MVS) Barnes et al. (2009) robustly and efficiently. Other works Yu et al. (2022b); Wang et al. (2022) propose utilizing monocular priors from pretrained models to achieve smooth and complete reconstruction. However, directly applying normal or depth priors Yu et al. (2022b) may lead to undesirable reconstruction results due to the unreliable nature of these priors. NeuRIS Wang et al. (2022) filters out unreliable priors based on the assumption that areas with rich 2D visual features are more error-prone. H2O-SDF Park et al. (2024) simply utilizes uncertainty prior to re-weight the normal prior loss. Nevertheless, both of these methods exhibit weak generalization capabilities. For example, the visual feature hypothesis Wang et al. (2022) struggles to differentiate between rich-textured planar areas. Similarly, prior-guided re-weighting Park et al. (2024) is further constrained by the domain gap between the prior model and the scene. On the other hand, DebSDF Xiao et al. (2024) effectively designs an uncertainty field based on a probabilistic model to guide the loss function, leading to more robust and accurate reconstructions. In contrast to methods with weaker generalization capabilities and those relying on probabilistic models, our proposed method, ND-SDF, learns a Normal Deflection field that enables the dynamic adaptation of priors based on their characteristics. This field directly models geometrically meaningful SO(3) residuals, resulting in highly detailed surface reconstructions while maintaining smoothness and robustness.
Figure 2: Overview of Our method. We utilize multi-resolution hash grids $\gamma_{L}$ as scene representation. The core of ND-SDF is the normal deflection field, which we represent using quaternions predicted by the deflection network (denoted as $f_{d}$). We align the deflected normals with the prior normals to learn the deviation between the scene and the priors. To distinctly supervise high and low-frequency areas, we employ an adaptive deflection angle prior loss, ensuring both smoothness and detail. Furthermore, we utilize the deflection angle $\Delta\theta$ to distinguish complex structures, enabling angle guided sampling and color loss to facilitate the recovery of intricate surface details. Lastly, we combine the unbiased rendering method Zhang et al. (2023) to ensure the generation of extremely thin structures indoors.
@@ 3 Method
The primary objective of our Normal Deflection Signed Distance Function (ND-SDF) is to reconstruct dense surfaces from calibrated multi-view images, with a primary focus on indoor scenes with established priors. To achieve this, we introduce a novel component known as the normal deflection field. This field is designed to quantify and learn the deviations between actual scene geometries and their corresponding normal priors. By integrating this field, our method can adaptively supervise both high and low-frequency regions in the scene. This adaptive supervision is crucial for preserving fine details while ensuring overall surface smoothness.
@@ 3.1 Preliminaries
Volume Rendering NeRF assumes that a ray $\textbf{r}(t)=\textbf{o}+t\textbf{v}$ is emitted from viewpoint o in direction v, where $t$ denotes the distance from the viewpoint. $N$ points are sampled along the ray, i.e. $\textbf{x}_{i}=\textbf{o}+t_{i}\textbf{v},i\in\{1,2...N\}$. Given each point’s volume density $\sigma_{i}$ and color $\textbf{c}_{i}$, the color of the ray $\hat{\textbf{C}}(\textbf{r})$ can be synthesized using volume rendering techniques as:
@@
where $\alpha_{i}$ is the opacity of $i$-th ray segment, $T_{i}$ denotes the transmittance of light reaching the $i$-th segment, $\delta_{i}=t_{i}-t_{i-1}$ denotes distance between adjacent samples.
A color loss supervised using ground-truth (GT) color is utilized to optimize the neural fields:
@@
where $\mathcal{R}$ is the sampled ray set, C(r) is the GT color.
SDF-induced Volume Rendering The rise of neural surface reconstruction is attributed to seminal works like VolSDF and Neus, which propose specific SDF to volume density conversion, allowing optimization via volume rendering. The final surface, denoted as $S$, is equivalent to the zero level-set of the implicit distance field, i.e., $S=\{\textbf{x}\in\mathbb{R}^{3}|s(\textbf{x})=0\}$, where $s(\textbf{x})$ is the SDF value.
We follow VolSDF Yariv et al. (2021), employing the Laplace Cumulative Distribution Function (CDF) to model the relationship between SDF and volume density:
@@
where $\Psi_{\beta}$ denotes the Laplace CDF, $\beta$ denotes the variance. Depth $\hat{D}(\textbf{r})$ and normal $\hat{\textbf{N}}(\textbf{r})$ at the surface intersection are synthesized using the same approach as that employed for synthesizing colors:
$\hat{D}(\textbf{r})=\sum_{i=1}^{N}T_{i}\alpha_{i}t_{i},\hat{\textbf{N}}(\textbf{r})=\sum_{i=1}^{N}T_{i}\alpha_{i}\textbf{n}_{i}.$ (4)
Here, $\textbf{n}_{i}$ is the analytical gradient of the SDF network at point $i$. Following MonoSDF Yu et al. (2022b), we utilize monocular depth and normal priors generated from pretrained model Eftekhar et al. (2021) to supervise the rendered depth and normal:
@@ 3.2 Normal deflection field
A significant challenge in indoor 3D reconstruction is achieving a balance between the smoothness of overall surfaces and the intricacy of complex structures. Traditional approaches relying solely on photometric loss have proven inadequate for accurately capturing smooth areas, particularly in texture-deficient regions. Later methods often leverage auxiliary data, such as normal priors, to enhance the reconstruction quality in textureless areas. However, the uniform application of normal priors across different scene types can impair the recovery of complex structures. This is primarily due to the variable reliability of these priors in diverse regions, where they may not accurately reflect the underlying geometry. To address this issue, we propose the development of a Normal Deflection Field. This field is designed to dynamically represent the existing 3D angular deviation between the actual scene normals and the provided normal priors. By doing so, it effectively circumvents potential misguidance arising from inconsistent priors, thereby enabling a more reliable and nuanced reconstruction of both smooth and complex indoor structures.
Specifically, we choose quaternion as the deflection form, which is a lightweight rotation representation. The quaternion can be parameterized by a single MLP $f_{d}$:
@@
where $\textbf{q}_{i}=(\textbf{q}_{i}^{0},\textbf{q}_{i}^{1},\textbf{q}_{i}^{2},\textbf{q}_{i}^{3})$ is the deflection quaternion (normalized default) for the sampled $\textbf{x}_{i}$. The quaternion at the surface intersection point is synthesized using volume rendering techniques, analogous to the way NeRF synthesizes colors:
@@ 3.3 Adaptive deflection angle prior loss
We assume that the learned deviation is encompassed within the actual deflected angle when the scene normal is deflected. This deflected angle ($\Delta\theta$) is computed as follows:
@@
where $\Delta\theta\in[0,\pi]$. Empirically, the deviation between the scene and priors increases as the structure becomes more complex. This intuitively suggests that the deflection angle is small in smooth regions and large in high-frequency zones. Based on this, the adaptive deflection angle normal prior loss is proposed to dynamically adjust the utilization of differing priors based on their characteristics:
$\begin{array}[]{cc}\mathcal{L}^{ad}_{normal}=&\sum_{\textbf{r}\in\mathcal{R}}{g^{d}(\Delta\theta)\mathcal{L}^{d}_{normal}(\hat{\textbf{N}}^{d}(\textbf{r}),{\textbf{N}}(\textbf{r}))}+g(\Delta\theta)\mathcal{L}_{normal}(\hat{\textbf{N}}(\textbf{r}),{\textbf{N}}(\textbf{r}))\end{array},$ (12)
where $g^{d}$ and $g$ are modulation functions that adjust the weights of the deflected and original normal loss terms based on the deflection angle. We similarly define an adaptive depth prior loss:
@@
As $\Delta\theta$ increases, $g^{d}(\Delta\theta)$ increases, and $g(\theta)$ decreases, since we should less apply priors in complex areas indicated by large deflection angles. The combined adaptive normal and depth prior loss is termed the adaptive deflection angle prior loss (as illustrated in Figure 2); see appendix for detailed definitions of the proposed modulation functions.
@@ 3.4 Deflection angle guided optimization
Through the utilization of the proposed adaptive prior loss, we dynamically adjust the utilization of various priors, significantly enhancing the quality of reconstruction. However, this alone is insufficient for generating more complex structures. Fundamentally, merely learning the deviation does not endow the capability to reconstruct additional details. Recognizing that large deflection angles indicate complex areas, we introduce three deflection angle guided optimization methods specifically designed to facilitate the recovery of more thin and fine-grained structures.
@@
Textureless areas, such as walls, constitute a significant portion of indoor scenes. These regions converge rapidly by leveraging prior knowledge, partly due to their inherent simplicity. To prevent continuous oversampling of well-learned smooth regions, we proactively sample more rays in complex areas to capture finer details. The sampling process is naturally guided by the learned deflection angles (see supplementary for details).
@@
To further enhance detail recovery, we impose additional photometric loss on complex areas identified by the deflection angles. The original color loss term is re-weighted based on the deflection angle, resulting in a re-weighted color loss.
@@
where ${\Delta\theta}(\textbf{r})$ denotes the actual deflected angle of the sampled ray r, and $w_{color}$ is the re-weighting function guided by ${\Delta\theta}$ (see appendix for details).
@@
Accurately reconstructing thin structures, such as chair legs, remains challenging. This limitation arises from the inherent bias issues in SDF-induced volume rendering, manifesting in two ways: (1) According to TUVR Zhang et al. (2023), the derivative of the rendering weight $\frac{\partial W(t)}{\partial t}$ is influenced by various angle differences between ray directions and scene normals. This results in a non-maximum value of $w$ at the surface intersection point, thereby degrading the surface quality; (2) For rays passing close to the object, the Laplace CDF assigns a high density to points near the surface, which leads to incorrect rendering weights, and subsequently affects the depth and normal compositing. This issue causes thin structures to disappear during training. Therefore, it is crucial to apply an unbiased rendering method during reconstruction. We illustrate this biased scenario in Fig. 3, highlighting the significance of ensuring unbiasedness.
Inspired by TUVR Zhang et al. (2023), we transform SDF to density using an unbiased function:
$\sigma(\textbf{r}({t_{i}}))=\frac{1}{\beta}\Psi_{\beta}\left(\frac{-f_{g}(\textbf{r}({t_{i}}))}{\left|f_{g}^{\prime}(\textbf{r}({t_{i}}))\right|}\right),$ (15)
where $\textbf{r}(t_{i})=\textbf{o}+t_{i}\textbf{v}$ represents a sampled point on the ray, and $f_{g}$ is the geometry network, which predicts the SDF value. This transformation effectively alleviates bias issues, as illustrated in Fig 3. However, we observe that simply applying it results in the overall surface failing to converge. Therefore, we only partially apply it to thin structures indicated by the learned deflection angles.
Figure 3: We present a scenario in which a ray passes close to a chair leg. Along the ray, point A is situated near the chair leg, while point B is located at the ray-ground intersection. In the absence of unbiasing, abnormal density and rendering weight are observed at point A, whereas density should be exclusively assigned to point B.